# E-Gen: Leveraging E-Graphs to Improve Continuous Representations of Symbolic Expressions

### v1

@@ C.2 Mistake Detection
Threshold Algorithm.
The threshold for mistake detection mentioned in Section 5.1 is computed using Algorithm 1. We first generate derivation steps based on each expression in the test set of the E-Gen corpus using SymPy, 80% of derivations are used for threshold calculation. For each derivation, we calculate the cosine similarities between consecutive steps. Then, the cosine similarities of steps with mistake are removed. We record the minimum cosine similarity of the remaining correct steps for each derivation. Finally, the average of the minimum cosine similarities is calculated as the threshold used for binary classification in the mistake detection task.
C.3 Formula Retrieval
@@
Formula retrieval is implemented as an extra task of semantic understanding. As discussed in Section 1, prior MIR studies rely heavily on contextual cues for semantic representation (Gao et al., 2017; Krstovski and Blei, 2018) rather than really explore mathematical property of expressions, resulting in suboptimal performance, especially in scenarios with limited surrounding text, such as textbook or mathematical derivations. To evaluate the model’s potential in this scenario, we design a test simulating pure formula retrieval based on the E-Gen corpus. Specifically, given a query expression, the top-$k$ most similar expressions are retrieved based on cosine similarity in the continuous embedding space.
@@
Table 14: Mathematical information retrieval accuracy (%) of seq2seq, CL Mean, and CL Max, compared against prior SemEmb model at different top-$k$ values. Accuracy denotes the proportion of top-$k$ ranked candidates that are semantically equivalent to the query expression.
The test set of E-Gen corpus which contains 8,077 expressions is tested, each expression is used as a query while the remaining expressions are used as candidates. Figure 4 shows an example in which top-$4$ most similar candidates to $\cos{x}$ are selected. As Table 14 shows, both seq2seq and contrastive learning models based on E-Gen corpus can effectively retrieve the most semantically relevant expressions from the candidate pool as $k$ increases. Conversely, due to a limited number of rewrites per expression, SemEmb struggles to identify semantically equivalent expressions when $k$ exceeds 10.
C.4 Embedding Algebra
Table 15 and 16 present additional examples from embedding algebra task. The seq2seq model correctly predicts most answers, outperforming other models. CL Mean slightly outperforms, CL Max, SemEmb, and GPT-4o. Notably, SemEmb and GPT-4o use a similar strategy to predict “$\boldsymbol{y}_{2}$”. For instance, Tests 6 to 11 evaluate the models’ understanding of function periodicity, where the seq2seq model trained on the E-Gen cluster-based corpus accurately predicts “$\boldsymbol{y}_{2}$” in most cases. In contrast, as discussed in Section 5.2, SemEmb, similar to GPT-4o, tends to imitate the transformation between “$\boldsymbol{x}_{1}$” and “$\boldsymbol{y}_{1}$” rather than correctly applying the periodicity of the function. In Test 6, “$\boldsymbol{x}_{1}:\sin(x)$” and “$\boldsymbol{y}_{1}:\sin(x+2\pi)$” are equivalent due to the $2\pi$ period of “$\sin(x)$”. However, since “$\boldsymbol{x}_{2}:\cot(x)$” has a period of $\pi$, the correct “$\boldsymbol{y}_{2}$” should be “$\cot(x+\pi)$”. The SemEmb model just simply imitates the transformation from “$\boldsymbol{x}_{1}$” to “$\boldsymbol{y}_{1}$” by adding “$2\pi$” to “$x$” in “$\cot(x)$”, and incorrectly converts “$\cot(...)$” into “$\csc(...)$”. This comparison highlights that the models trained with E-Gen corpus shows a better understanding in mathematical rules and transformations, rather than relying on superficial "looks-like" predictions.
Figure 5: Two mistake detection conversation examples in GPT-4o prompt. The mistakes in the derivation are in red. In the left one, GPT-4o successfully detects the mistake step, but also reports a false positive mistake step. In the right one, GPT-4o fails to detect the mistake in the derivation at all.
@@ C.4 Embedding Algebra
Table 15: Additional examples of embedding algebra with seq2seq and CL Mean models. Both models are trained using E-Gen corpus and this evaluation shows the difference between seq2seq and CL Mean models. $\hat{\boldsymbol{y}}_{2}$ is the prediction from the corresponding model, and $\boldsymbol{y}_{gt}$ is the ground truth. The expressions in red are wrong predictions. The seq2seq model offers superior performance over the CL Mean model.
@@
Table 16: Additional examples of embedding algebra comparing the embedding model (CL Max) trained on the E-Gen corpus presented in this manuscript with prior work (SemEmb) and GPT-4o. $\hat{\boldsymbol{y}}_{2}$ is the prediction from the corresponding model, and $\boldsymbol{y}_{gt}$ is the ground truth. The expressions in red are wrong predictions.

### v2

@@ 5.1 Evaluation Tasks
For this task, step-by-step mathematical derivations are generated for each expression in the E-Gen corpus, and mistakes are introduced into randomly selected steps within these derivations. The goal is to ask the models to identify erroneous steps by classifying transformations between consecutive steps as either “mistake” or “no mistake”. To achieve this, a semantic similarity threshold for binary classification is calculated using only derivations generated from expressions in the training set of the E-Gen corpus, as described in Algorithm 1. Specifically, cosine similarities are computed between consecutive steps in each derivation, excluding steps with mistakes. The minimum cosine similarity among the correct steps is recorded for each derivation, and the average of these minimum values is used as the threshold. A step is flagged as a mistake if its cosine similarity with the preceding step falls below this threshold.
Derivation seq2seq GPT-4o
$-\frac{1}{8}(\sinh(\cosh^{-1}(7-\frac{x}{-8})))^{-1}$
$-\frac{1}{8}(\sinh(\cosh^{-1}(7+\frac{x}{8})))^{-1}$
$-\frac{1}{8}(\sqrt{6+\frac{x}{8}}\sqrt{8+\frac{x}{8}})^{-1}$bigstar
$-\frac{1}{8}(6+\frac{x}{8})^{-\frac{1}{2}}(8+\frac{x}{8})^{-\frac{1}{2}}$
$-\frac{1}{8}(6+\frac{x}{8})^{-\frac{1}{2}}(8-7x)^{-\frac{1}{2}}$
$5/\csc(\csc^{-1}(1/\ln(\frac{x}{-5})))$
$5/\csc(\csc^{-1}(1/\ln(\frac{1}{-5}x)))$
$5/{\color[rgb]{1,0,0}\sec}(\csc^{-1}((\ln(\frac{1}{-5}x))^{-1}))$bigstar
$5/(1-(\ln(\frac{1}{-5}x))^{2})^{-\frac{1}{2}}$
$5\sqrt{1-(\ln(\frac{1}{-5}x))^{2}}$
Examples of mistake detection are shown in Table 7. Some mistakes are particularly challenging to identify, as they may closely resemble the structure of the preceding step and appear deceptively correct. Conversely, steps with significant syntactic changes that are mathematically equivalent can be misclassified as errors. These challenges highlight the importance of robust semantic understanding to ensure accurate mistake detection.
Model Precision Recall F1
no mistake seq2seq 96.40 94.69 95.54
CL Mean 97.59 92.52 94.99
CL Max 97.78 91.93 89.14
SemEmb 92.92 83.49 87.95
mistake seq2seq 74.68 81.61 77.99
CL Mean 69.33 88.10 77.60
CL Max 67.96 89.14 77.12
SemEmb 44.46 67.50 53.61
The test set comprises 18,462 derivation steps, of which 2,974 steps contain mistakes, with the remainder being error-free. As shown in Table 8, precision, recall, and F1-score are used to evaluate model performance. Models trained on E-Gen corpus demonstrate strong effectiveness in identifying potential mistakes and better generalizability on this OOD task, significantly outperforming the SemEmb approach.
Embedding Algebra.
Embedding algebra is a classic task to evaluate if embeddings capture semantic information of a word/token. Techniques such as word2vec (Mikolov, 2013) and GloVe (Pennington et al., 2014) exhibit the ability to perform analogy-based reasoning through algebraic operations on their representation vectors, enabling solutions to analogies like “Berlin is to Germany as Paris is to France”. Extending this task to mathematical expressions allow us to assess whether models truly understand mathematical transformations or merely rely on surface-level structural similarity. For a given triplet of expressions $x_{1}$, $y_{1}$, and $x_{2}$, we compute:
$f(\hat{y}_{2})=-f(x_{1})+f(y_{1})+f(x_{2})$ (2)
where $f$ denotes a function $f:\mathcal{X}\rightarrow\mathcal{Z}$, which maps an expression $x$ to its representation vector $z$ in the latent space. The expression whose embedding vector has the highest cosine similarity to $f(\hat{y}_{2})$ is selected as the predicted answer, excluding the original expressions $x_{1}$, $y_{1}$, and $x_{2}$ from consideration.
For this experiment, 584 analogy examples are manually constructed. The entire E-Gen corpus serves as the search pool, with expressions equivalent to $x_{2}$ and $y_{2}$ removed to ensure uniqueness of the correct answer. Any necessary expressions are added to complete the analogy.
Model Accuracy (%)
seq2seq 70.38
CL Mean 64.73
CL Max 50.34
SemEmb 54.85
GPT-4o 39.60
As shown in Table 10, the seq2seq model achieves the highest accuracy, indicating its ability to learn underlying mathematical rules and handle basic substitutions effectively. In contrast, SemEmb exhibit low accuracy on this task and have a tendency to imitate the transformation between $x_{1}$ and $y_{1}$, leading to “look-alike” predictions rather than true understanding of the mathematical rules, as seen in Test 1, 4, 5 in Table 9.
5.2 Comparison with GPT-4o
To further assess the quality of vector representations of symbolic expressions, a comparative analysis is conducted between the seq2seq model, trained on the E-Gen corpus, and the state-of-the-art large language model GPT-4o (Achiam et al., 2023) on two tasks.
Mistake Detection.
For mistake detection, a small test set is randomly sampled from the mistake detection test set described earlier. This subset consists of 322 derivation steps, including 50 steps containing mistakes, while the remainder are error-free. GPT-4o is first provided with an example derivation in the prompt and asked to identify potential mistakes. If it fails to detect the mistake, explicit feedback indicating the erroneous step is provided until it correctly understands the task. Once verified, GPT-4o is tested on queries from the sampled test set.
Model Precision Recall F1
no mistake seq2seq 96.98 94.49 95.72
CL Mean 97.64 91.18 94.30
CL Max 97.78 91.93 94.77
GPT-4o 93.58 90.84 92.19
mistake seq2seq 73.68 84.00 78.50
CL Mean 64.71 88.00 74.58
CL Max 67.96 89.14 77.12
GPT-4o 56.90 66.00 61.11
As shown in Table 11, our approach outperforms GPT-4o across all evaluation metrics. While GPT-4o demonstrates a recall of 66%, suggesting its capability of detecting a notable portion of errors, its low precision indicates a tendency to misclassify correct transformations as mistakes, leading to a high rate of false positives.
Table 7 illustrates examples of GPT-4o’s misjudgments, which commonly occur in complex transformations, particularly those involving function operators or relatively intricate arithmetic computations. For instance, in Example 1, GPT-4o fails to recognize the equivalence between syntactically different expressions (e.g., $\sinh(\cosh^{-1}(x))=\sqrt{x+1}\sqrt{x-1}$). In Example 2, it incorrectly classifies an operator substitution in step 3 as a valid derivation, failing to detect the mistake.
Embedding Algebra.
We also conduct the embedding algebra task on GPT-4o, which performs worse than both the seq2seq and CL models, achieving only 39.60% accuracy on the test set as shown in Table 10. Table 9 provides examples comparing the performance of the seq2seq model and GPT-4o on embedding algebra tests. Similar to the mistake detection test, an example query is provided to GPT-4o to verify its understanding before proceeding with the test.
While the seq2seq model consistently makes accurate predictions by adhering to mathematical rules, GPT-4o demonstrates only a partial understanding of certain mathematical properties. For instance, it correctly predicts the periodicity of “$\csc(x)$” but often mimics the structure of “$y_{1}$” rather than applying the underlying mathematical transformations. In Test 5, GPT-4o incorrectly subtracts “$\pi/2$” from “$x$” in “$\csc(x)$”, imitating the structure of “$y_{1}$”, while the correct answer should be “$\csc(x+\pi/2)$”, which is mathematically equivalent to “$\csc(x)$”. A similar error occurs in Test 8, where GPT-4o predicts “$y_{2}$” as “$\operatorname{csch}^{-1}(x)$” by replicating “$y_{1}:\csc^{-1}(x)$” instead of applying the correct mathematical transformation. This comparison highlights the effectiveness of E-Gen corpus in helping models to build up understanding in mathematical rules and transformations.
6 Conclusion
In this work, we enhance semantic representations of symbolic expressions by developing E-Gen, a novel mathematical corpus generation framework based on the e-graph data structure. E-Gen shows strong capability in the generation of a scalable, cluster-based corpus with high diversity in mathematical transformation. To evaluate its effectiveness, two training approaches based on seq2seq and contrastive learning respectively are implemented to capture equivalence relation between expressions. Our experimental results demonstrate the efficacy of these embedding models across a variety of downstream tasks, including clustering, semantic understanding beyond syntactic similarity, mistake detection, and mathematical analogies. Notably, these semantic representations outperform prior approach and GPT-4o in both quantitative and qualitative tests. This work provides an algorithmic foundation for processing symbolic mathematics, and the vector-based representations can easily integrate with vector embeddings of other data modalities.
Limitations
This work introduces and evaluates the potential of a novel mathematical corpus generation framework to augment mathematical semantic understanding. While the results are promising, there are several areas for improvement.
First, extending the dataset to include a wider range of mathematical operators used in published datasets, such as ArXMLiv (Kohlhase et al., 2024) and ARQMath (Mansouri et al., 2022b), would improve the applicability of the learned embeddings to real-world mathematical data.
Following that, more efficient grammar enumeration techniques, such as those based on the Earley algorithm (Earley, 1970), could facilitate the extension of the E-Gen corpus to support more complex operators and higher-arity expressions. In this manuscript, equivalent expressions are generated by enumerating the grammar obtained from the e-graph using a simple recursive function. Additionally, incorporating variable characteristics, such as dimensionality, phase, and/or bounds, into the embeddings remains an unexplored but promising direction for improving their expressiveness.
Another limitation is the scarcity of real-world datasets that focus on symbolic mathematics. While tasks like mathematical information retrieval have been explored in competitions such as NTCIR (Zanibbi et al., 2016a) and ARQMath (Mansouri et al., 2022b), these typically involve a mixture of symbolic mathematics and natural language in queries and results. Integration of our embedding methods with natural language vectors remains an open research challenge with significant potential for advancing mathematical information retrieval and reasoning.
Acknowledgments
We would like to thank the University of Illinois for its support in facilitating this research. We would also like to extend our gratitude to the National Center for Supercomputing Applications (NCSA) for providing access to high-performance computing resources.
Appendix A Cluster-Based Corpus
A.1 Cluster Example
Initial expression: $(\sin(2x+5))^{-8}$
$(\sin(5-(-2)\sin(\sin^{-1}(x))))^{-8}$
$(\csc(\pi+(-5-2x)))^{8}$
$(\cos(\pi/2-2x-5)))^{-8}$
$\left(\frac{\cot(2x+5)}{\cos(2x+5)}\right)^{8}$
$(\sin(2(x+\pi)+5))^{-8}$
Initial expression: $\sec^{-1}(7x+6)/(-7)$
$\cos^{-1}(\frac{1}{6-(-7x)})/(-7)$
$(\sec^{-1}(-7x-6)-\pi))\times\frac{1}{7}$
$(\frac{\pi}{2}-\csc^{-1}(7x+6))/(-7)$
$(-1)(\sec^{-1}(7/\sec(\cos^{-1}(x))+6))/7$
$(\cos^{-1}(\cot(\tan^{-1}(7x+6))))/(-7)$
Initial expression: $(\csc(x/7))^{-3}-(-3)$
$31/\left(\csc\left(\frac{\tan(\tan^{-1}x)}{7}\right)\right)^{3}+$ 3
$(\tan(\cot^{-1}(\sin(\frac{x}{7}))))^{-3}+3$
$3+\left(1/\cos(\sec^{-1}(\sin(\frac{x}{7})))\right)^{3}$
$-1(-3-(\csc(\frac{x}{7}))^{-3})$
$(\sin(\frac{x}{7}))^{3}+3$
$(\sin(\csc^{-1}(\sin(7^{-1}x)))^{-3}-(-3)$
Initial expression: $\frac{d}{dx}(-4x-8)^{-4}+7x$
$(-1)\times(8-(-4x))^{-5}\times 16+7$
$-16\times(4x+8)^{-5}+7$
$7+(-16)\times\left(-1/(-4x-8)^{5}\right)$
$7-(-16)\times(-1)\times(8-(-4x))^{-5}$
$0-(-16(-4x-8)^{-5}-7)$
Initial expression: $\frac{d}{dx}(-9\sinh(-7x+2))$
$63/\operatorname{sech}(-7x+2)$
$-9/(-1/7\cosh(7x-2))$
$63\cosh(2-7x)$
$\frac{63}{2}(\exp(2-7x)+\exp(7x-2))$
$7\times 9/\operatorname{sech}(7x-2)$
Table 12: Additional examples of equivalent expressions clusters generated with E-Gen. Expressions listed below each initial expression are all equivalent to it.
As discussed in Section 3.2, the key distinction of the E-Gen corpus from prior datasets is its cluster-based organization of equivalent expressions, rather than equivalent expression pairs. Table 12 presents additional examples of these clusters. Each cluster consists of an initial expression along with numerous equivalent rewrites. This diverse set of transformations significantly enhances pretrained models’ ability to understand and generalize mathematical semantics, leading to improved performance in downstream tasks.
A.2 Generation Efficiency
Efficiency is a key property of any corpus generation scheme. To assess the computational performance of E-Gen, we measure the average time required for e-graph saturation and expression rewrite extraction. On an Intel Xeon E5-2666 CPU, the e-graph saturation process takes 36.83ms on average, and the extraction takes an average of 1.46s per expression. These evaluations are conducted with a 25-token length limit and a 600s timeout per initial expression.
Appendix B Training Details
Configuration Value
Transformer Architecture
Model Dimension 512
Attention Heads 8
Feedforward Dimension 2048
Encoder Layers 6
Decoder Layers 6
Dropout -
Optimizer
Optimizer AdamW
Learning Rate $1\times 10^{-4}$
Weight Decay $1\times 10^{-2}$
Scheduler
Scheduler CosineAnnealWarmRestarts
$T_{0}$ 10
$T_{\text{mult}}$ 2
$\eta_{\text{min}}$ $1\times 10^{-8}$
Training Parameters
Label Smoothing 0.1
Batch Size 256
Epochs 20
Gradient Clipping 4.0
Hardware Configuration
CPU AMD EPYC 7763 (64-Core)
GPU NVIDIA L40S (46GB)
Table 13: Hyperparameters and hardware specifications for training the seq2seq, CL Mean, and CL Max models. The transformer-based models are optimized using AdamW with a cosine annealing warm restart scheduler. Training was conducted on an AMD EPYC 7763 CPU and a Nvidia L40S GPU.
Both the seq2seq and contrastive learning models are trained using a transformer architecture, optimized with the AdamW optimizer (Loshchilov, 2017), and scheduled with CosineAnnealingWarmRestarts scheduler (Loshchilov and Hutter, 2016). The specific training parameters and time are detailed in Table 13 and Table 14 respectively.
Model Training Time (h:min)
seq2seq 38:04
CL Mean 31:56
CL Max 29:59
Table 14: Training time of seq2seq, CL Mean, and CL Max models.
Appendix C Experiments
C.1 K-Means Clustering
Accuracy Calculation.
$\textit{acc}=\frac{1}{K}\sum_{i=1}^{K}\textit{acc}_{i}$ (3)
C.2 Formula Retrieval
@@
Formula retrieval is implemented as an additional task to assess the models’ semantic understanding of mathematical expressions. As discussed in Section 1, prior MIR studies heavily rely on contextual cues for semantic representation (Gao et al., 2017; Krstovski and Blei, 2018) rather than directly capturing the intrinsic mathematical property of expressions. This limitation results in suboptimal performance, particularly in scenarios with limited surrounding text, such as textbook or mathematical derivations. To evaluate the models’ performance in this scenario, we design a pure formula retrieval task using the E-Gen corpus. Given a query expression, the top-$k$ most similar expressions are retrieved based on cosine similarity in the latent space.
We evaluate formula retrieval using the E-Gen test set, which contains 8,077 expressions. Each expression serves as a query, while the remaining expressions act as retrieval candidates. Figure 4 illustrates an example where the top-$4$ most similar candidates to $\cos{x}$ are retrieved in the latent space. As shown in Table 15, both seq2seq and CL models trained on E-Gen corpus effectively retrieve semantically relevant expressions from the candidate pool as $k$ increases. Conversely, SemEmb, which is constrained by a limited number of equivalent rewrites per expression, struggles to identify semantically equivalent expressions when $k$ exceeds 10.
@@
Table 15: Formula retrieval accuracy (%) of seq2seq, CL Mean, and CL Max, compared against prior SemEmb model at different top-$k$ values. Accuracy denotes the proportion of top-$k$ ranked candidates that are semantically equivalent to the query expression.
C.3 Embedding Algebra
Table 16 and 17 provide additional examples from embedding algebra task. The seq2seq model correctly predicts most of the answers, outperforming other models. CL Mean slightly outperforms, CL Max, SemEmb, and GPT-4o. Notably, SemEmb and GPT-4o use a similar strategy to predict “$\boldsymbol{y}_{2}$”. For instance, Tests 6 to 11 specifically evaluate the models’ understanding of function periodicity, where the seq2seq model trained on the E-Gen corpus accurately predicts “$\boldsymbol{y}_{2}$” in most cases. In contrast, as discussed in Section 5.2, SemEmb, similar to GPT-4o, tends to imitate the transformation between “$\boldsymbol{x}_{1}$” and “$\boldsymbol{y}_{1}$” rather than correctly applying the periodicity of the function. For instance, in Test 6, “$\boldsymbol{x}_{1}:\sin(x)$” and “$\boldsymbol{y}_{1}:\sin(x+2\pi)$” are equivalent due to the $2\pi$ period of “$\sin(x)$”. However, since “$\boldsymbol{x}_{2}:\cot(x)$” has a period of $\pi$, the correct “$\boldsymbol{y}_{2}$” should be “$\cot(x+\pi)$”. The SemEmb model just simply imitates the transformation from “$\boldsymbol{x}_{1}$” to “$\boldsymbol{y}_{1}$” by adding “$2\pi$” to “$x$” in “$\cot(x)$”, and incorrectly converts “$\cot(\ldots)$” into “$\csc(\ldots)$”. This comparison highlights that the models trained with E-Gen corpus shows a better understanding of mathematical rules and transformations, rather than relying on superficial "looks-like" predictions.
Figure 5: Two examples of mistake detection conversations in the GPT-4o prompt. Errors in the derivation are highlighted in red. In the left example, GPT-4o correctly identifies the erroneous step but also incorrectly flags a false positive. In the right example, GPT-4o fails to detect the mistake entirely.
@@ C.3 Embedding Algebra
Table 16: Additional examples from the embedding algebra evaluation comparing the seq2seq and CL Mean models, both trained on the E-Gen corpus. $\hat{\boldsymbol{y}}_{2}$ represents model’s prediction, while $\boldsymbol{y}_{gt}$ denotes the ground truth. Incorrect predictions are highlighted in red. The seq2seq model offers superior performance over the CL Mean model.
@@
Table 17: Additional examples from the embedding algebra evaluation comparing the CL Max model trained on the E-Gen corpus with prior work SemEmb and GPT-4o. $\hat{\boldsymbol{y}}_{2}$ represents model’s prediction, while $\boldsymbol{y}_{gt}$ denotes the ground truth. Incorrect predictions are highlighted in red. All three models have comparable performance.